# Asymmetric thin-shell wormholes in the Kalb-Ramond background: Observational characteristics and extra photon rings

### v2

@@ Abstract
In this paper, we utilize the ray-tracing method to conduct an in-depth study of the observational images of asymmetric thin-shell wormholes in the Kalb-Ramond field. Initially, we calculate the null geodesics and effective potential energy of ATSW, and investigate the variations in the photon sphere radius and critical impact parameter under different values of charge $Q$ and Lorentz-violating parameter $l$. Based on these calculations, we determine the photon deflection angles and trajectories within this space-time structure. Specifically, depending on the photon impact parameters, the photon trajectories can be categorized into three types. By using a thin accretion disk as the sole background light source and incorporating two classical observational radiation models, we find that under conditions of equal mass $M$, charge quantity $Q$, and Lorentz-violation parameter $l$, the asymmetric thin-shell wormhole exhibits unique observational features such as additional lensing rings and photon ring clusters. Furthermore, distinct from black holes, as the charge quantity $Q$ and the Lorentz-violation parameter $l$ increase, the coverage area of the specific additional halo also expands correspondingly.
@@ I Introduction
In 2015, the Laser Interferometer Gravitational-Wave Observatory (LIGO) made the first detection of gravitational waves, directly confirming the existence of binary stellar-mass black holes-a landmark achievement in astrophysics LIGOScientific:2016sjg . Subsequently, in 2019, the Event Horizon Telescope (EHT) released images of the supermassive black holes at the centers of the M87 galaxy and Sagittarius A*  EventHorizonTelescope:2019dse , providing further evidence for the existence of such supermassive black holes. The black hole images display a central dark region surrounded by a bright ring: the dark area results from the black hole’s intense gravity bending and capturing light, preventing it from reaching distant observers, thus forming what is known as the black hole shadow. The outer bright ring is referred to as the photon ring (or critical curve) Gralla:2019xty .
The study of black hole shadows has always been a hot topic in the field of physics Perlick:2021aok ; Zhang:2024lsf ; Zeng:2025olq ; Hou:2022gge ; Wang:2025ihg ; Zhang:2025vyx ; He:2024amh ; Fan:2024rsa . Synge was the first to calculate and determine the boundary of the black hole shadow , while Luminet, pioneered the study of the angular radius equation for the photon capture region in a Schwarzschild black hole Synge:1966okc ; Luminet:1979nyg . In recent years, this field has seen remarkable developments within the framework of modified gravity theories. On the observational front, Heino Falcke and his colleagues were the first to propose the concept of the black hole shadow, and their work directly contributed to the development and advancement of EHT Falcke:1999pj ; Zeng:2025kqw . Although the EHT’s observations are largely consistent with the black hole shadow predictions of General Relativity, they do not rule out the existence of certain Ultra-Compact Objects (UCOs) Bambi:2008jg ; Banerjee:2019nnj ; Abdikamalov:2019ztb . In fact, objects such as wormholes and boson stars could potentially cast shadows very similar to those of black holes Tsukamoto:2021fpp ; Wielgus:2020uqz ; Wang:2023nwd ; Rosa:2023qcv ; Rosa:2022tfv ; Cunha:2016bjh . Consequently, finding methods to effectively distinguish UCOs from black holes has become a crucial research direction.
@@
The structure of this paper is as follows. In Section II, an ATSW is constructed within the KR field. We investigate its geodesics and effective potential, and subsequently calculate the deflection and trajectories of photons in this ATSW space-time. Section III.1 examines the transfer functions and observational appearances of both a BH and an ATSW, each surrounded by a thin accretion disk and under the same mass parameter, for two emission models. Finally, our conclusions are presented in Section IV.
@@ II Null Geodesics and Effective Potential of the ATSW in the Kalb-Ramond Field
In the theoretical framework, $R$ denotes the Ricci scalar, $\eta$ and $\xi_{1}$, $\xi_{2}$ are coupling constants representing the non-minimal coupling terms between gravity and the KR field, while $\mathcal{L}_{\mathrm{N}}$ refers to the Lagrangian of the electromagnetic field. In our previous work, we investigated the influence of the charge $Q$ and the Lorentz-violating parameter $l$ on black hole shadows Tan:2025pya . Specifically, the study focused on a KR black hole surrounded by a thin accretion disk. It was found that an increase in the charge $Q$ and the Lorentz-violating parameter $l$ leads to a reduction in the size of the black hole shadow, as well as in the direct image, lensing ring, and photon ring. The present research shifts focus to the optical characteristics of an Asymmetric Thin-Shell Wormhole within the KR framework. This investigation is motivated by two key aspects: First, previous studies on Bardeen-type ATSWs have shown that the magnetic charge $g$ significantly affects their optical appearance He:2024yeg . This intriguing result prompts us to examine whether the charge $Q$ and the Lorentz-violating parameter $l$ similarly induce changes in the optical appearance of the KR field ATSW, analogous to the case of the KR black hole Zeng:2024ptv . Second, we aim to determine whether shadow-related characteristics can provide a viable method for distinguishing an ATSW from a black hole.
The ATSW consists of two space-times with distinct mass parameters, $M_{1}$ and $M_{2}$, connected by a thin shell (ie, the throat), with the entire ATSW manifold denoted by $M=M_{1}\cup M_{2}$. Based on Eq. (1), with the cosmological constant $\Lambda$ set to zero, we derive the following equation Duan:2023gng
@@
Here, $M$, $Q$, and $l$ represent the black hole’s mass, charge, and Lorentz-violating parameter, respectively. Expanding Eq.(4) as a Taylor series yields
@@
In which, $i=1,2$ denotes the two space-time regions, $R$ represents the position of the thin shell and satisfies the condition $R>\max\left\{r_{h_{1}},r_{h_{2}}\right\}$, where $r_{h_{1}}$ and $r_{h_{2}}$ represent the event horizon radii of black holes with mass parameters $M_{1}$ and $M_{2}$, respectively. Considering that gravity acts as the sole interaction when photons pass through the throat, the 4-momentum $p^{A}(A=t,r,\theta,\phi)$ of the photon remains conserved.
Since the space-time is symmetric, photons possess two conserved quantities: $E_{i}=-p_{t_{i}}$ (energy) and $L_{i}=p_{\phi_{i}}$ (angular momentum)-when propagating along geodesics, and their motion satisfies the geodesic equations
$\frac{\left(p_{i}^{r_{i}}\right)^{2}}{F_{i}\left(r_{i}\right)}+\frac{p_{\varphi_{i}}^{2}}{r_{i}^{2}}=\frac{p_{t_{i}}^{2}}{F_{i}\left(r_{i}\right)},$ (6)
where $p_{i}^{r_{i}}=\mathrm{d}x^{r_{i}}/\mathrm{d}\lambda$ represents the four-momentum of the photon in the space-time, and $\lambda$ denotes the affine parameter. This affine parameter can be redefined as $\lambda=\lambda/E$, and $b$ is the impact parameter, which can be expressed as $b=|L|/E$. From the above relations, we derive the following
@@
For the unstable circular orbit of a photon, the photon sphere must satisfy the conditions
$V_{e\!f\!f_{i}}\left(r_{ph_{i}}\right)=\frac{1}{b_{c_{i}}^{2}},\quad V_{i}^{\prime}\left(r_{ph_{i}}\right)=0.$ (9)
Here, $b_{c}$ and $r_{p}$ denote the critical impact parameter and the photon sphere radius, respectively. The primary objective of this paper is to investigate a method for distinguishing an ATSW from a black hole within the KR field, specifically under the condition where the event horizon is obscured by the photon sphere. To this end, we assume an observer is situated in the space-time with mass parameter $M_{1}=1$, and the mass parameter of the other space-time $M_{2}$ is set to $M_{2}=k$. Under this configuration, the parameters $k$ and the throat radius $R$ must satisfy the following conditions Wang:2020emr ; Peng:2021osd
@@
In essence, within the valid range of the parameter $k$, its specific value does not qualitatively affect the subsequent results and discussion. Therefore, we adopt the approach used in prior studies and set the value of $k=1.2$ Peng:2021osd ; Guo:2022iiy . Consequently, the relationship between the impact parameters $b_{1}$ and $b_{2}$ for the space-times $M_{1}$ and $M_{2}$ can be derived as follows
@@
To study the effects of the charge $Q$ and the Lorentz-violating parameter $l$ on the photon sphere radius $r_{p}$ and the critical collision parameter $b_{c}$, calculations were conducted using the formula for the photon sphere radius with $M_{1}=1$ and $M_{2}=1.2$. The results are shown in Table 1 and Table 2. Based on the relevant research, the charge $Q$ is limited to the range from 0 to $0.6$, and the Lorentz-violating parameter $l$ is limited to the range from 0 to $0.1$. It was observed that when $Q=0$ and $l=0$, the scenario corresponds to the Schwarzschild ATSW, and when $l=0$, it corresponds to the Reissner-Nordström case.
@@
Table 1: The variation of photon sphere radius $r_{p_{i}}$ and critical impact parameter $b_{c_{i}}$ with respect to charge $Q$, in KR spacetime with Lorentz violating parameter $l=0.01$. The charge values are taken as $Q=$ 0, 0.1, 0.3, 0.50,\,0.1,\,0.3,\,0.5. The discussion covers two spacetime configurations with mass parameters $M_{1}=1$ and $M_{2}=1.2$ respectively.
@@
Table 2: The variation of the photon sphere radius $r_{p_{i}}$ and the critical impact parameter $b_{c_{i}}$ with respect to the Lorentz-violating parameter $l$, in the KR spacetime with ATSW (charge $Q=0.3$). The values of the Lorentz-violating parameter $l$ are taken as $l=$ 0, 0.01, 0.05, 0.10,\,0.01,\,0.05,\,0.1. The analysis is conducted for two distinct spacetime configurations with mass parameters $M_{1}=1$ and $M_{2}=1.2$, respectively.
@@
(a) The effective potential $V_{e\!f\!f}$ of BH in the KR field under varying Lorentz-violating parameter $l$.
(b) The effective potential $V_{e\!f\!f}$ energy of ATSW in the KR field varies with the change of the Lorentz-violating parameter $l$.
@@
The effective potential energies of BH and ATSW in the KR field are depicted in Fig.  1  and  2 , respectively. By examining the effective potential energy of BH (with charge $Q=0.3$) in Fig. 1 (a) and  2 (a), we discern three distinct photon trajectories as illustrated in the diagram on the lower right side. For the scenario where $l=0.01$ and $Q=0.1$, as shown by the black line in Fig. 1 , when $b_{1}>b_{c_{1}}=5.11841$, the gravitational influence results in the deflection of the photon trajectory. When $b_{1}=b_{c_{1}}$, the photon will perpetually orbit around the BH in a circular path. When $b_{1}<b_{c_{1}}$, due to the intense gravitational force, the photon will directly plunge into the BH. Furthermore, it is noted that with the increase in charge $Q$ and the Lorentz-violating parameter $l$, the effective potential energy of the BH also rises, and the position corresponding to the peak shifts inward.
In Fig. 1 (b) and Fig. 2 (b), we plot the effective potential of the ATSW as a function of the Lorentz-violating parameter $l$ for a fixed charge $Q=0.1$. The dashed line on the left corresponds to spacetime $M_{2}$, while the solid line on the right corresponds to spacetime $M_{1}$. The effective potential of spacetime $M_{2}$ is constrained by a factor of $Z^{2}$. From the observations of Fig. 1 (b) and Fig. 2 (b), we note that, for a given charge $Q$ or Lorentz-violating parameter $l$, photons exhibit three distinct trajectories Gralla:2019xty . For the case where $l=0.1$ and $Q=0.1$, as indicated by the black line in Fig. 1 (b): When $b_{1}<Zb_{c_{1}}=3.708$, the photon falls from spacetime $M_{1}$ into spacetime $M_{2}$ and moves toward infinity.When $Zb_{c_{2}}<b_{1}<b_{c_{1}}=5.11841$, the photon enters the potential barrier in spacetime $M_{2}$ and then returns to spacetime $M_{1}$. When $b_{1}<Zb_{c_{2}}$, the photon moves to the potential barrier in spacetime $M_{1}$ and then travels toward infinity in $M_{1}$.
@@
$\frac{1}{b_{i}^{2}}-\frac{f_{i}\left(r_{i}\right)}{r_{i}^{2}}=\frac{1}{r_{i}^{4}}\left(\frac{\mathrm{~d}r_{i}}{\mathrm{~d}\varphi}\right)^{2}.$ (12)
@@
$G_{i}\left(x_{i}\right)=\left(\frac{\mathrm{d}x_{i}}{\mathrm{~d}\varphi}\right)^{2}=\frac{1}{b_{i}^{2}}-x_{i}^{2}\left[\left(1+{Q^{2}x_{i}^{2}}-{2M_{i}x_{i}}\right)+\left(1+{2Q^{2}x_{i}^{2}}\right)l+\left(1+{3Q^{2}x_{i}^{2}}\right)l^{2}\right],$ (13)
@@
$\varphi_{1}\left(b_{1}\right)=2\int_{0}^{x_{1}^{\min}}\frac{\mathrm{d}x_{1}}{\sqrt{G_{1}\left(x_{1}\right)}},\quad b_{1}>b_{c_{1}},$ (14)
@@
$\varphi_{1}^{*}\left(b_{1}\right)=\int_{0}^{1/R}\frac{dx_{1}}{\sqrt{G_{1}\left(x_{1}\right)}},\quad b_{1}<b_{c_{1}},$ (15)

### v3

@@ Abstract
In this paper, we utilize the ray-tracing method to conduct an in-depth study of the observational images of asymmetric thin-shell wormholes in the Kalb-Ramond field. Initially, we calculate the null geodesics and effective potential energy of the asymmetric thin-shell wormhole, and investigate the variations in the photon sphere radius and critical impact parameter under different values of charge $Q$ and Lorentz-violation parameter $l$. Based on these calculations, we determine the photon deflection angles and trajectories within this space-time structure. Specifically, depending on the photon impact parameters, the photon trajectories can be categorized into three types. By using a thin accretion disk as the sole background light source and incorporating two classical observational radiation models, we find that under conditions of equal mass $M$, charge quantity $Q$, and Lorentz-violation parameter $l$, the asymmetric thin-shell wormholes exhibit unique observational features such as additional lensing rings and photon ring clusters. Furthermore, distinct from black holes, as the charge quantity $Q$ and the Lorentz-violation parameter $l$ increase, the coverage area of the specific additional halo also expands correspondingly.
@@ I Introduction
In 2015, the Laser Interferometer Gravitational-Wave Observatory (LIGO) made the first detection of gravitational waves, directly confirming the existence of binary stellar-mass black holes-a landmark achievement in astrophysics LIGOScientific:2016sjg . Subsequently, in 2019, the Event Horizon Telescope (EHT) released images of the supermassive black holes at the centers of the M87 galaxy and Sgr A EventHorizonTelescope:2019dse , providing further evidence for the existence of such supermassive black holes. The black hole images display a central dark region surrounded by a bright ring: the dark area results from the black hole’s intense gravity bending and capturing light, preventing it from reaching distant observers, thus forming what is known as the black hole shadow. Within the shadow, the bright ring contains a very thin, embedded ring—commonly referred to as the photon ring—which theoretically corresponds to the so-called "critical curve" Gralla:2019xty .
The study of black hole shadows has always been a hot topic in the field of physics Perlick:2021aok ; Zhang:2024lsf ; Zeng:2025olq ; Hou:2022gge ; Wang:2025ihg ; Zhang:2025vyx ; He:2024amh ; Fan:2024rsa ; Zeng:2025kqw . Synge was the first to calculate and determine the boundary of the black hole shadow, while Luminet, pioneered the study of the angular radius equation for the photon capture region in a Schwarzschild black hole Synge:1966okc ; Luminet:1979nyg . Falcke demonstrated that the shadow of Sgr A* could be observed via very long-baseline interferometry at sub-millimeter wavelengths, under the assumption that the accretion flow is optically thin in this spectral region Falcke:1999pj . Additionally, extensive research has also been conducted on the signature of modified theory of gravity on the photon sphere and shadow, and some meaningful progress has been achieved Guo:2018kis ; Wang:2018prk . Although the EHT’s observations are largely consistent with the black hole shadow predictions of General Relativity, they do not rule out the existence of certain Ultra-Compact Objects (UCOs) Bambi:2008jg ; Banerjee:2019nnj ; Abdikamalov:2019ztb . In fact, objects such as wormholes and boson stars could potentially cast shadows very similar to those of black holes Tsukamoto:2021fpp ; Wielgus:2020uqz ; Wang:2023nwd ; Rosa:2023qcv ; Rosa:2022tfv ; Cunha:2016bjh . Consequently, finding methods to effectively distinguish UCOs from black holes has become a crucial research direction.
@@
The structure of this paper is as follows. In Section II, an ATSW is constructed within the KR field. We investigate its geodesics and effective potential, and subsequently calculate the deflection and trajectories of photons in this ATSW space-time. Section III examines the transfer functions and observational appearances of both a BH and an ATSW, each surrounded by a thin accretion disk and under the same mass parameter, for two emission models. Finally, our conclusions are presented in Section IV.
@@ II Null Geodesics and Effective Potential of the ATSW in the Kalb-Ramond Field
In the theoretical framework, $R$ denotes the Ricci scalar, $\eta$ and $\xi_{1}$, $\xi_{2}$ are coupling constants representing the non-minimal coupling terms between gravity and the KR field, while $\mathcal{L}_{\mathrm{N}}$ refers to the Lagrangian of the electromagnetic field. In our previous work, we investigated the influence of the charge $Q$ and the Lorentz-violation parameter $l$ on black hole shadows Tan:2025pya . Specifically, the study focused on a KR black hole surrounded by a thin accretion disk. It was found that an increase in the charge $Q$ and the Lorentz-violation parameter $l$ leads to a reduction in the size of the black hole shadow, as well as in the direct image, lensing ring, and photon ring. The present research shifts focus to the optical characteristics of an Asymmetric Thin-Shell Wormhole within the KR framework. This investigation is motivated by two key aspects: First, previous studies on Bardeen-type ATSWs have shown that the magnetic charge $g$ significantly affects their optical appearance He:2024yeg . This intriguing result prompts us to examine whether the charge $Q$ and the Lorentz-violation parameter $l$ similarly induce changes in the optical appearance of the KR field ATSW, analogous to the case of the KR black hole Zeng:2024ptv . Second, we aim to determine whether shadow-related characteristics can provide a viable method for distinguishing an ATSW from a black hole.
The ATSW consists of two space-times with distinct mass parameters, $M_{1}$ and $M_{2}$, connected by a thin shell (ie, the throat), with the entire ATSW manifold denoted by $M=M_{1}\cup M_{2}$. Based on Eq. (1), with the cosmological constant $\Lambda$ set to zero, we can derive the following equation Duan:2023gng
@@
Here, $M$, $Q$, and $l$ represent the black hole’s mass, charge, and Lorentz-violation parameter, respectively. Expanding Eq. (4) as a Taylor series in small $l$ yields
@@
In which, $i=1,2$ denotes the two space-time regions, $R$ represents the position of the thin shell and satisfies the condition $R>\max\left\{r_{h_{1}},r_{h_{2}}\right\}$, where $r_{h_{1}}$ and $r_{h_{2}}$ represent the event horizon radii of black holes with mass parameters $M_{1}$ and $M_{2}$, respectively. Considering that gravity is the sole interaction as photons traverse the throat, the energy $p_{t}$ and axial angular momentum $p_{\phi}$ are all conserved.
Since the space-time is symmetric, we choose the plane $\theta=\pi/2$ (passing through the coordinate origin) for analysis; photons then possess two conserved quantities: $E_{i}=-p_{t_{i}}$ (energy) and $L_{i}=p_{\phi_{i}}$ (angular momentum) when propagating along geodesics, and their motion satisfies the geodesic equations
$\frac{\left(p_{i}^{r_{i}}\right)^{2}}{F_{i}\left(r_{i}\right)}+\frac{p_{\phi_{i}}^{2}}{r_{i}^{2}}=\frac{p_{t_{i}}^{2}}{F_{i}\left(r_{i}\right)},$ (6)
where $p_{i}^{r_{i}}=\mathrm{d}x^{r_{i}}/\mathrm{d}\lambda$ represents the four-momentum of the photon in the space-time, and $\lambda$ represents the affine parameter. It is conventional to redefine the affine parameter using the conserved energy as $\lambda=\lambda/E$, which allows the impact parameter $b$ — a dimensionless quantity characterizing the photon’s trajectory — to be precisely defined as the ratio of the conserved angular momentum to the conserved energy: $b=|L|/E$ Luo:2023wru ; He:2024yeg . Based on these relations, we derive the following expression
@@
For a photon to move on a circular orbit, the orbit must satisfy the conditions for an extremum of the effective potential
$V_{e\!f\!f_{i}}\left(r_{ph_{i}}\right)=\frac{1}{b_{c_{i}}^{2}},\quad V_{e\!f\!f_{i}}^{\prime}\left(r_{ph_{i}}\right)=0,\quad\frac{d^{2}V_{e\!f\!f_{i}}(r_{ph_{i}})}{dr^{2}}\Big|_{r=r_{ph_{i}}}<0.$ (9)
Here, $b_{c_{i}}$ and $r_{ph_{i}}$ denote the critical impact parameter and the photon sphere radius, respectively. When the condition $b=b_{c}$ is satisfied — meaning the light ray’s impact parameter is infinitely close to the radius of the photon sphere — the ray will orbit the black hole infinitely many times.
The stability of this orbit is determined by the sign of the second derivative. An unstable circular orbit, which defines the photon sphere, corresponds to a local maximum. By combining Eqs. (5) and  (9), we obtain the expression for $r_{ph_{i}}$ and $b_{c_{i}}$
$r_{ph_{i}}=\frac{3M_{i}+\sqrt{9M_{i}^{2}-4\mathcal{AA}\left(2Q^{2}+4Q^{2}l+6Q^{2}l^{2}\right)}}{2\mathcal{AA}},$ (10)
$b_{c_{i}}=\frac{\sqrt{-27M_{i}^{4}+36M_{i}^{2}Q^{2}\mathcal{BB}-8Q^{4}\mathcal{BB}^{2}-9M_{i}^{3}\sqrt{9M_{i}^{2}-8O^{2}\mathcal{BB}}+8M_{i}Q^{2}\mathcal{BB}\sqrt{9M_{i}^{2}-8Q^{2}\mathcal{BB}}}}{\sqrt{2}\sqrt{\mathcal{AA}\left(-M_{i}^{2}+Q^{2}\mathcal{BB}\right)}},$ (11)
where
$\mathcal{AA}=1+l+l^{2},\mathcal{BB}=1+3l+6l^{2}+5l^{3}+3l^{4}.$ (12)
This leads to the expression for the area $S_{p_{i}}$ enclosed by the photon sphere radius in the two-dimensional plane
$S_{p_{i}}=\pi r_{ph_{i}}^{2}.$ (13)
The primary objective of this paper is to investigate a method for distinguishing an ATSW from a black hole within the KR field, specifically under the condition where the event horizon is enclosed by the photon sphere. To this end, we assume an observer is situated in the space-time with mass parameter $M_{1}=1$, and the mass parameter of the other space-time $M_{2}$ is set to $M_{2}=k$. Under this configuration, the parameters $k$ and the throat radius $R$ must satisfy the following conditions Wang:2020emr ; Peng:2021osd
@@
In essence, within the valid range of the parameter $k$, its specific value does not qualitatively affect the subsequent results and discussion. Therefore, we adopt the approach used in prior studies and set the value of $k=1.2$ Peng:2021osd ; Guo:2022iiy . The local tetrad frame in the neighborhood of the thin-shell wormhole in spacetime $M_{1,2}$ can be expressed as
$\mathbf{e}_{t_{i}}^{a}\equiv F_{i}^{-\frac{1}{2}}(R)\left(\frac{\partial}{\partial t_{i}}\right)^{a},\quad\mathbf{e}_{r_{i}}^{a}\equiv\sqrt{F_{i}}(R)\left(\frac{\partial}{\partial r_{i}}\right)^{a},$ (15)
which are then related by the Lorentz transformation Langlois:2001uq
$\begin{pmatrix}\mathbf{e}_{t_{2}}^{a}\\ \mathbf{e}_{r_{2}}^{a}\end{pmatrix}=\Lambda(\theta_{1}-\theta_{2})\begin{pmatrix}\mathbf{e}_{t_{1}}^{a}\\ \mathbf{e}_{r_{1}}^{a}\end{pmatrix}$. (16)
We define $\Lambda(\theta_{i})$ and $\theta_{i}$ within the range as follows
$\Lambda(\theta_{i})$ $=\begin{pmatrix}\cosh\theta_{i}&\sinh\theta_{i}\\ \sinh\theta_{i}&\cosh\theta_{i}\end{pmatrix},\quad\theta_{i}=\sinh^{-1}\left(\frac{\epsilon_{i}\dot{R}}{\sqrt{F_{i}(R)}}\right).$ (17)
Where $\dot{R}$ denotes the velocity of the moving thin shell. By definition, the metric of spacetime $M_{i}$ is continuous (i.e., $\mathbf{g}_{ab}^{M_{1}}(R)=\mathbf{g}_{ab}^{M_{2}}(R)$) Nakao:2013hba ; Wang:2020emr . Consequently, the two conserved quantities $E_{i}=-p_{t_{i}}$ and $L_{i}=p_{\phi_{i}}$ along the geodesic are expressed as
$-\frac{E_{2}}{\sqrt{F_{2}(R)}}$ $=-\frac{\cosh(\theta_{1}-\theta_{2})}{\sqrt{F_{1}(R)}}E_{1}+\frac{\sinh(\theta_{1}-\theta_{2})}{\sqrt{F_{1}(R)}}p_{R}^{r_{1}},$ $L_{1}$ $=L_{2}.$ (18)
Where $p_{R}^{r_{1}}$ denotes the value of $p^{r_{1}}$ on the thin shell. By further combining Eqs. (7) and  (18), the relationship between the impact parameters $b_{1}$ and $b_{2}$ for spacetime $M_{1}$ and spacetime $M_{2}$ can be derived as follows
@@
To study the effects of the charge $Q$ and the Lorentz-violation parameter $l$ on the photon sphere radius $r_{ph_{i}}$ and the critical collision parameter $b_{c_{i}}$, calculations were conducted using the formula for the photon sphere radius with $M_{1}=1$ and $M_{2}=1.2$. The results are shown in Table. (1) and  (2). And the variations of the photon sphere radius and its corresponding area with respect to the charge $Q$ and the Lorentz-violation parameter $l$ are plotted in Figs. 1 and  2 . Based on the relevant research, the charge $Q$ is limited to the range from 0 to $0.5$, and the Lorentz-violation parameter $l$ is limited to the range from 0 to $0.1$. It was observed that when $Q=0$ and $l=0$, the scenario corresponds to the Schwarzschild ATSW, and when $l=0$, it corresponds to the Reissner-Nordström case. From Figs. 1 , 2  and Table. (1), (2), it is observed that the impact parameter $b_{c_{i}}$, photon sphere radius $r_{p_{i}}$, and the corresponding enclosed area $S_{p_{i}}$ decrease with increasing charge $Q$ and Lorentz-violation parameter $l$.
@@
Table 1: The variation of photon sphere radius $r_{p_{i}}$ and critical impact parameter $b_{c_{i}}$ with respect to charge $Q$, in KR spacetime with Lorentz violation parameter $l=0.01$. The charge values are taken as $Q=$ 0, 0.1, 0.3, 0.50,\,0.1,\,0.3,\,0.5. The discussion covers two spacetime configurations with mass parameters $M_{1}=1$ and $M_{2}=1.2$ respectively.
@@
Table 2: The variation of the photon sphere radius $r_{p_{i}}$ and the critical impact parameter $b_{c_{i}}$ with respect to the Lorentz-violation parameter $l$, in the KR spacetime with ATSW (charge $Q=0.3$). The values of the Lorentz-violation parameter $l$ are taken as $l=$ 0, 0.01, 0.05, 0.10,\,0.01,\,0.05,\,0.1. The analysis is conducted for two distinct spacetime configurations with mass parameters $M_{1}=1$ and $M_{2}=1.2$, respectively.
Figure 1: (color online) In spacetime $M_{1}$, panels (a), (b), and (c) respectively depict the variation of the photon sphere radius $r_{p_{1}}$, impact parameter $b_{c_{1}}$, and the enclosed area $S_{p_{1}}$ of the photon sphere with the Lorentz-violation parameter $l$.
Figure 2: (color online) In spacetime $M_{2}$, the variation of the photon sphere radius $r_{p_{2}}$, impact parameter $b_{c_{2}}$, and the area $S_{p_{2}}$ enclosed by the photon sphere with the Lorentz-violation parameter $l$, is depicted in panels (a), (b), and (c), respectively.
@@
(a) The effective potential $V_{e\!f\!f}$ of BH in the KR field under varying Lorentz-violation parameter $l$.
(b) The effective potential $V_{e\!f\!f}$ energy of ATSW in the KR field varies with the change of the Lorentz-violation parameter $l$.
@@
The effective potential energies of BH and ATSW in the KR field are depicted in Figs.  3  and  4 , respectively. By examining the effective potential energy of BH (with charge $Q=0.3$) in Figs. 3 (a) and  4 (a), we discern three distinct photon trajectories as illustrated in the diagram on the lower right side. For the scenario where $l=0.01$ and $Q=0.1$, as shown by the black line in Fig. 3 , when $b_{1}>b_{c_{1}}=5.11841$, the gravitational influence results in the deflection of the photon trajectory. When $b_{1}=b_{c_{1}}$, the photon will perpetually orbit around the BH in a circular path. When $b_{1}<b_{c_{1}}$, due to the intense gravitational force, the photon will directly plunge into the BH. Furthermore, it is noted that with the increase in charge $Q$ and the Lorentz-violation parameter $l$, the effective potential energy of the BH also rises, and the position corresponding to the peak shifts inward.
In Figs. 3 (b) and  4 (b), we plot the effective potential of the ATSW as a function of the Lorentz-violation parameter $l$ for a fixed charge $Q=0.1$. The dashed line on the left corresponds to spacetime $M_{2}$, while the solid line on the right corresponds to spacetime $M_{1}$. The effective potential of spacetime $M_{2}$ is constrained by a factor of $Z^{2}$. From the observations of Figs. 3 (b)  4 (b), we note that, for a given charge $Q$ or Lorentz-violation parameter $l$, photons exhibit three distinct trajectories Gralla:2019xty . For the case where $l=0.1$ and $Q=0.1$, as indicated by the black line in Fig. 3 (b): When $b_{1}<Zb_{c_{1}}=3.708$, the photon falls from spacetime $M_{1}$ into spacetime $M_{2}$ and moves toward infinity.When $Zb_{c_{2}}<b_{1}<b_{c_{1}}=5.11841$, the photon enters the potential barrier in spacetime $M_{2}$ and then returns to spacetime $M_{1}$. When $b_{1}<Zb_{c_{2}}$, the photon moves to the potential barrier in spacetime $M_{1}$ and then travels toward infinity in $M_{1}$.
@@
$\frac{1}{b_{i}^{2}}-\frac{f_{i}\left(r_{i}\right)}{r_{i}^{2}}=\frac{1}{r_{i}^{4}}\left(\frac{\mathrm{~d}r_{i}}{\mathrm{~d}\phi}\right)^{2}.$ (20)
@@
$G_{i}\left(x_{i}\right)=\left(\frac{\mathrm{d}x_{i}}{\mathrm{~d}\varphi}\right)^{2}=\frac{1}{b_{i}^{2}}-x_{i}^{2}\left[\left(1+{Q^{2}x_{i}^{2}}-{2M_{i}x_{i}}\right)+\left(1+{2Q^{2}x_{i}^{2}}\right)l+\left(1+{3Q^{2}x_{i}^{2}}\right)l^{2}\right].$ (21)
@@
$\varphi_{1}\left(b_{1}\right)=2\int_{0}^{x_{1}^{\min}}\frac{\mathrm{d}x_{1}}{\sqrt{G_{1}\left(x_{1}\right)}},\quad b_{1}>b_{c_{1}}.$ (22)
@@
$\varphi_{1}^{*}\left(b_{1}\right)=\int_{0}^{1/R}\frac{dx_{1}}{\sqrt{G_{1}\left(x_{1}\right)}},\quad b_{1}<b_{c_{1}}.$ (23)